OpenStudy (anonymous):

Find the standard form of the equation of the ellipse with the given characteristics. vertices: (- 2, 2), (- 2, 16) minor axis of length 8

OpenStudy (anonymous):

ANSWER CHOICES: (x-2)^2/16+(y+9)^2/49=1 (x+2)^2/16+(y-9)^2/49=1 (x+2)^2/49+(y-9)^2/16=1 (x-9)^2/49+(y+2)^2/16=1 (x+9)^2/49+(y-2)^2/16=1

OpenStudy (mertsj):

The length of the major axis is 2a. The length of the minor axis is 2b. The center is at the midpoint of the major axis.

OpenStudy (mertsj):

Use that information to find a, b and the center (h,k) and plug into this equation: \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (jdoe0001):

notice the vertices of the ellipse (-2, 2) and (-2, 16) notice the x-coordinate doesn't change, but the y-coordinate does that means the ellipse major axis is the y-axis, or is "vertical" keep in mind that center of the ellipse, is half-way between the vertices half-way between (-2, 2) and (-2, 16), let's use the midpoint formula for kicks :) \(\bf \left(\cfrac{(-2)+(-2)}{2}, \cfrac{(2)+(16)}{2}\right) \implies (-2, 9)\) so the center is at (-2, 9) the distance from the center to either vertex is "a" over the "y" is moved from (-2, 2 ) to (-2, 9), that is 7 units thus a = 7 the minor axis is given, that is, 8 units, thus b = 8 so (h, k) = (-2, 9) # center a = 7 # major axis unit b = 8 # minor axis uniit see above as Mertsj showed, the equation for the ellipse, and put your values in :)